OpenStudy (anonymous):

Hey , How would I normally know the integration is from this point to the other point , I'll post the question as an example. in a sec

OpenStudy (anonymous):

find the given area: the area below the x axis and above y=x^2 , x=2 and the x-axis how would I find which points ? for example in this one it's from 0 to 2 , I already know 2 but what about 0 ? is there a normal way to find out ? \[\int\limits_{0}^{2} ...\]

ganeshie8 (ganeshie8):

the area \(\color{red}{above}\) the x axis and \(\color{red}{below}\) y=x^2 , x=2 and the x-axis

ganeshie8 (ganeshie8):

right ?

ganeshie8 (ganeshie8):

first step for problems like these is to "sketch" the curves and get a feel for the region that u wanto find the area

OpenStudy (anonymous):

No , sorry but it says like I've posted :D maybe the prof was wrong I'll check the book later , but can you please answer my question

ganeshie8 (ganeshie8):